# Breaking the Curse of Dimensionality: On the Stability of Modern Vector Retrieval

### v1

@@ 6 Stability of Filtered Vector Search
We now move to the second major vector-based retrieval setting we analyze in this paper, namely filtered similarity search. As in the previous section, we will begin by defining the formal mathematical model for filtered vector search we consider and then introduce our main theorem. In summary, we make the following contributions in this section:
  1. 1.
    We present a formal mathematical model of penalty-based filtered vector search, which captures several common practical instantiations of this problem.
  2. 2.
    We prove that filtered near neighbor search in the penalty-based model is stable if the distance penalty function for a filter mismatch is set sufficiently high.
  3. 3.
    We demonstrate, both theoretically and experimentally, that filtered search with sufficiently high penalties can be stable even if the underlying standard near neighbor search problem is unstable.
@@ 6.1 Theoretical Model
  • $\mathcal{Q}=\{(q_{1},A_{q_{1}}),(q_{2},A_{q_{2}}),\dots\}$ is a set of query tuples where each $q_{i}\in\mathbb{R}^{m}$ and each $A_{q_{i}}\subseteq A$ is a set of query attributes.
@@
This definition of filtered near neighbor search places no restriction on how the filter function $f$ operates. For instance, if the query and document attributes are discrete, then a natural definition of $f$ might be to output 1 if $A_{q}\subseteq A_{d}$ and 0 otherwise. Furthermore, our definition extends to other filtered search settings. In the approximate window search problem, (Engels et al., 2024), each document is associated with a numeric label, and each query aims to find the nearest documents within arbitrary label ranges. In this case, we can model $A_{q}\subset\mathbb{R}$ as a continuous range of real numbers and $A_{d}\in\mathbb{R}$ as a scalar value. The filter function $f$ in this setting would output 1 if $A_{d}\in A_{q}$ and 0 otherwise.
@@
We note that this formal model of fused distance functions captures nearly all filtered vector search algorithms in the literature. Some approaches, such as NHQ (Wang et al., 2023) and CAPS (Gupta et al., 2023) set the penalty value via heuristics or parameter tuning. On the other hand, many approaches in the literature, such as Acorn (Patel et al., 2024), Filtered DiskANN (Gollapudi et al., 2023), and window search Engels et al. (2024) never consider points that violate the query filter predicate throughout the entire search process. These algorithms can also be modeled by our framework by setting $\alpha$ to be infinite. Thus, we emphasize that our formalization of filtered vector search closely aligns with nearly all algorithmic approaches in the literature.
@@ 6.2 Theoretical Results
We now present our main theorem of this section, which proves that we can achieve stability in the filtered search setting by choosing a sufficiently large distance penalty value. In this theoretical analysis, we assume that the attributes are drawn from some distribution such that the filter mismatch probability is at most some constant $p_{\max}$ where $0<p_{\max}<1$.
@@
We defer the full proof of this theorem to Appendix B.
Remark 6.4.
Theorem 6.3 holds regardless of whether the underlying vector search problem (without constraints) is stable or unstable since our proof handles both cases directly. The latter case is particularly interesting as it implies that, under the assumptions of Theorem 6.3, the addition of filters can transform an unstable vector search problem into a stable one.
Remark 6.5.
This model of filtered vector search is quite general and can encompass an arbitrary number of conjunctions and disjunctions of constraints simply by defining the attribute sets $A_{q}$ and $A_{d}$ accordingly. We only require that the constant $p_{\max}$ remains non-trivial.
@@ 6.3 Experiments
In this section, we empirically validate Theorem 6.3 on a family synthetic vector search problems with 10,000 documents and 100 queries and varying dimensionality. We unit normalize all vectors so that the constant $\Delta\geq 2$ and set the parameter $p_{\max}$ to be $0.5$ which satisfies the conditions of the theorem. Thus, as predicted by our theory, a sufficient condition for stability based on these parameters is setting the penalty value $\alpha$ to be greater than $8$. For this experiment, we start with a base vector search problem that is unstable and we consider three different choices for the penalty value: in the first case we take the penalty to be 8.1, namely slightly above the minimum threshold for stability; secondly, we set the penalty value to be smaller than the threshold suggested by Theorem 6.3 and have it decay with the dimensionality; finally, we consider the case of a penalty value of 0. As suggested by the proof of Theorem 6.3, we want to engineer filter mismatches such that there is a negative correlation between the distance function and attribute matches. We simulate this phenomenon by setting filter matches and mismatches based on each document’s average distance to all queries; documents that are close to queries on average become filter mismatches while documents further away are matches.
@@ 7 Stability of Sparse Vector Search
We now turn our attention to the final vector search setting we study in this work: sparse embedding retrieval. We make the following contributions in this section:
  1. 1.
    We introduce a mathematical model for sparse vector retrieval, including a novel definition of the concentration of importance property first informally described by (Bruch et al., 2024).
  2. 2.
    We prove that concentration of importance, coupled with a new property we call overlap of importance and an additional mild assumption, implies sparse vector search stability.
  3. 3.
    We validate our theorem empirically, showing that concentration of importance and overlap of importance together provide stability, but one property without the other remains insufficient.
@@ 7.1 Theoretical Model
[…]
@@ 7.2 Theoretical Results
The above theorem is, to our knowledge, the first to prove a sufficient condition for stability in the context of sparse vector search, which is a particularly interesting setting given the high dimensionalities of learned sparse embeddings, such as SPLADE, used in practice.
We first point out that we state the theorem using $\liminf\mathrm{RelVar}_{m}$ instead of $\lim\mathrm{RelVar}_{m}$. This is convenient because bounding $\liminf$ away from zero is sufficient to prove stability without needing to prove that the relative variance terms converge, which may or may not happen.
The theorem shows that sparse vector search is stable whenever $X=(2-2\gamma)^{1/p}$ is less than $Y=\frac{\rho 2^{1/p}}{1-\pi}\left(\alpha^{1/p}-(1-\alpha)^{1/p}\right)$. These two constants represent different geometric configurations for the vectors $q$ and $d$. The quantity $X$ is an upper bound on the expected distance when $q$ and $d$ overlap substantially in their head coordinates. In this case, the vectors partially align and the distance shrinks. In contrast, $Y$ represents the expected distance in the opposite configuration where $T_{q}$ and $T_{d}$ are disjoint. Here the vectors place their largest coordinates in opposite directions, and hence the distance grows.
Equality $X=Y$ corresponds to a single algebraic relationship between parameters $\alpha,\gamma$ and $\pi$ where the two geometric configurations happen to produce the exact same distance. Geometrically, this means that placing mass on overlapping head coordinates reduces the distance by precisely the same amount that shifting the mass into disjoint coordinates would increase it. This cancellation requires the parameters $(\alpha,\gamma,\pi)$ to satisfy a rigid relation and therefore occurs on a set of measure zero on the parameter space. Everywhere else for all practically relevant parameterizations, the two configurations remain geometrically distinct, which ensures that the distribution of the distances does not collapse in the limit, which we require for stability.
7.3 Empirical Validation of Theorem Assumptions
We empirically validate the assumptions of Theorem 7.4 on $30,522$-dimensional SPLADE-V3 Lassance et al. (2024) embeddings across several benchmark datasets. We verify concentration of importance, overlap of importance, and finally check the gap condition $Y>X$ required by the theorem.
Concentration of importance.
Overlap of importance.
For random query-document pairs we compute the overlap statistic
$S=\sum_{i\in T_{q}\cap T_{d}}\min\{q_{i}^{p},d_{i}^{p}\}$
Corpus $\alpha$ $\kappa$ $R$ $\rho$ $\gamma$ $\hat{\pi}$ $d_{\max}/d_{\min}$ YXY-X
HotpotQA 0.85 24 3 0.9167 0.00157 0.5194 3.6966 0.0291
TREC-COVID 0.85 24 6 0.9443 0.00236 0.8213 4.7285 2.5830
Natural-Questions 0.85 18 6 0.9537 0.00137 0.5335 2.9501 0.1324
MS MARCO 0.95 18 8 0.9691 0.00049 0.3267 3.2981 0.1159
NFCorpus 0.85 18 8 0.9615 0.0012 0.5412 2.3999 0.1713
@@ 7.4 Stability of Synthetic Sparse Embeddings
Figure 5 illustrates how we create two regimes in our datasets, one that satisfies both concentration and overlap of importance and the other that only satisfies concentration. Furthermore, we also construct two more regimes, one where we only have overlap of important coordinates and the last one where we do not fulfill either condition. Figure 6 shows the results of the experiments when we set $\alpha=0.83$, $\gamma=0.2$, and $\pi=0.5$ for the overlap probability. In particular, the CoI-only regime highlights that even strongly-concentrated embeddings exhibit distance collapse as the dimension grows when there is not enough overlap in their heads. These results confirm that neither concentration nor overlap of importance are sufficient conditions for stability on their own.
qqshared headddEmbedding dimensionsqqhead(q)(q)ddhead(d)(d)Embedding dimensions
Figure 5: Two semantic-locality regimes in the synthetic dataset generator. Left: query $q$ and document $d$ share a semantic head, so their top-$\kappa$ coordinates overlap (CoI + overlap). Right: $q$ and $d$ are each concentrated (CoI) but their heads are disjoint, illustrating the CoI-only regime.
@@ 8 Conclusion
Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under Grant No. 2313998, 1845763, 2316235, and 2403237. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation.
Appendix A Stability of Multi-Vector Search
@@
Theorem 5.5 A multi-vector search instance is stable if and only if
$\lim_{m\to\infty}\frac{\mathrm{Var}\left[\mathrm{Agg}(Q,X)\right]}{\mathbb{E}\left[\mathrm{Agg}(Q,X)\right]^{2}}\neq 0$
@@
This result follows directly from the prior proofs of (Beyer et al., 1999) and (Durrant & Kabán, 2009) since the original proofs require only that the distance metric is non-negative. We provide a full, self-contained argument here for completeness.
As a matter of notation, let $X_{m}=\mathrm{Agg}(Q,D)$ be the random variable for the aggregated distance in an $m$-dimensional space. Let its expected value be $\mathbb{E}[X_{m}]$ and its variance be $\text{Var}[X_{m}]$. The database consists of $n$ samples of this random variable, $\{X_{m,1},\dots,X_{m,n}\}$. Let $\mathrm{DMIN}_{m}=\min_{i}X_{m,i}$ and $\mathrm{DMAX}_{m}=\max_{i}X_{m,i}$.
$\implies$ This direction follows the argument from (Beyer et al., 1999). We will show that if $\lim_{m\to\infty}\frac{\text{Var}[X_{m}]}{\mathbb{E}[X_{m}]^{2}}=0$, then the multi-vector search problem is unstable. Define a normalized random variable $Y_{m}=\frac{X_{m}}{\mathbb{E}[X_{m}]}$. Its expectation is $\mathbb{E}[Y_{m}]=1$ and its variance is $\text{Var}[Y_{m}]=\frac{\text{Var}[X_{m}]}{\mathbb{E}[X_{m}]^{2}}$ The condition is thus equivalent to $\lim_{m\to\infty}\text{Var}[Y_{m}]=0$. Since $\lim_{m\to\infty}\mathbb{E}[Y_{m}]=1$ and $\lim_{m\to\infty}\text{Var}[Y_{m}]=0$, the random variable $Y_{m}$ converges in probability to 1. This is denoted $Y_{m}\xrightarrow{p}1$. Let $\{Y_{m,1},\dots,Y_{m,n}\}$ be the normalized distances for the $n$ items in the database. Since $\min()$ and $\max()$ are continuous functions, Slutsky’s Theorem implies that $\min(Y_{m,i})\xrightarrow{p}1$ and $\max(Y_{m,i})\xrightarrow{p}1$. The ratio $\frac{\mathrm{DMAX}_{m}}{\mathrm{DMIN}_{m}}=\frac{\max_{i}Y_{m,i}}{\min_{i}Y_{m,i}}$.Since both the numerator and denominator converge in probability to 1, their ratio also converges in probability to 1. By definition, $\frac{\mathrm{DMAX}_{m}}{\mathrm{DMIN}_{m}}\xrightarrow{p}1$ means that for any $\epsilon>0$, we have $\lim_{m\to\infty}P\left(\left|\frac{\mathrm{DMAX}_{m}}{\mathrm{DMIN}_{m}}-1\right|\leq\epsilon\right)=1$. Since $\mathrm{DMAX}_{m}\geq\mathrm{DMIN}_{m}$, this is equivalent to $\lim_{m\to\infty}P(\mathrm{DMAX}_{m}\leq(1+\epsilon)\mathrm{DMIN}_{m})=1$. This is precisely the definition of an unstable search problem.
@@
$\frac{\mathrm{DMIN}_{m}}{\mathrm{DMAX}_{m}}\leq\frac{X_{m,i}}{\mathbb{E}[X_{m}]}\leq\frac{\mathrm{DMAX}_{m}}{\mathrm{DMIN}_{m}}$
From the instability assumption, we know both the lower bound $(\frac{\mathrm{DMIN}_{m}}{\mathrm{DMAX}_{m}})$ and the upper bound $(\frac{\mathrm{DMAX}_{m}}{\mathrm{DMIN}_{m}})$ converge in probability to 1 By the Squeeze Theorem, the random variable $Y_{m}$ must also converge in probability to 1. A random variable that converges in probability to a constant must have a variance that converges to zero. Therefore:
$\lim_{m\to\infty}\text{Var}[Y_{m}]=\lim_{m\to\infty}\frac{\text{Var}[X_{m}]}{\mathbb{E}[X_{m}]^{2}}=0$
@@
$\frac{\mathrm{DMAX}^{\prime}}{\mathrm{DMIN}^{\prime}}=\frac{\max_{k}\min_{d\in D_{k}}\delta(q,d)}{\mathrm{DMIN}}\geq\frac{1}{c}\frac{\mathrm{DMAX}}{\mathrm{DMIN}}>1.$
where the last inequality follows from strong stability. Thus, since $\frac{\mathrm{DMAX}^{\prime}}{\mathrm{DMIN}^{\prime}}>1$ almost surely, it follows that $\frac{\mathrm{DMAX}^{\prime}}{\mathrm{DMIN}^{\prime}}\not\to_{p}1$, completing the proof. ∎
@@
Lemma A.1 (Dan’s Favorite Inequality (Spielman, 2018)).
@@
where the last two inequalities follow from applying Lemma A.1 and Lemma 5.8 respectively. This completes the proof.
Remark A.2.
Stability follows from applying Dan’s Favorite Inequality (Lemma A.1) to reduce the general multi-vector search problem to multi-vector search over singleton query sets, namely the setting in Lemma 5.8, at the cost of an $\frac{1}{k}$ overhead factor. The $1/k$ factor suggests that stability weakens as the query set size grows, which is also consistent with empirical observations.
Remark A.3.
A natural question to ask based on the proof of Theorem 5.9 is where we utilize the fact that our aggregation function is the Chamfer distance. In fact, the properties of Chamfer distance are critical to the argument, but the application is subtle. In particular, we rely on the minimum operator of the Chamfer distance in Lemma 5.8 to connect the stability of the induced single vector search problem to the vector set setting. Since we apply Lemma 5.8 to prove Theorem 5.9, our main theorem in turn also crucially relies on the Chamfer distance.
Appendix B Stability of Filtered Vector Search

### v2

@@ 5 Stability of Filtered Vector Search
We now move to the second major vector-based retrieval setting we analyze in this paper, namely filtered similarity search. As in the previous section, we will begin by defining the formal mathematical model for filtered vector search we consider and then introduce our main theorem.
@@ 5.1 Theoretical Model
  • $\mathcal{Q}=\{(q_{1},A_{q_{1}}),(q_{2},A_{q_{2}}),\dots\}$ is a set of query tuples where each $q_{i}\in\mathbb{R}^{m}$ and each $A_{q_{i}}\subseteq A_{Q}$ is a set of query attributes.
@@
This definition of filtered near neighbor search places no restriction on how the filter function $f$ operates. For instance, if the query and document attributes are discrete, then a natural definition of $f$ might be to output 1 if $A_{q}\subseteq A_{d}$ and 0 otherwise. Furthermore, our definition extends to other filtered search settings. In the approximate window search problem (Engels et al., 2024), each document is associated with a numeric label, and each query aims to find the nearest documents within arbitrary label ranges. In this case, we can model $A_{q}\subset\mathbb{R}$ as a continuous range of real numbers and $A_{d}\in\mathbb{R}$ as a scalar value. The filter function $f$ in this setting would output 1 if $A_{d}\in A_{q}$ and 0 otherwise.
@@
We note that this formal model of fused distance functions captures nearly all filtered vector search algorithms in the literature. Some approaches, such as NHQ (Wang et al., 2023) and CAPS (Gupta et al., 2023) set the penalty value via heuristics or parameter tuning. On the other hand, many approaches in the literature, such as Acorn (Patel et al., 2024), Filtered DiskANN (Gollapudi et al., 2023), and window search (Engels et al., 2024) never consider points that violate the query filter predicate throughout the entire search process. These algorithms can also be modeled by our framework by setting $\alpha$ to be infinite. Thus, we emphasize that our formalization of filtered vector search closely aligns with nearly all algorithmic approaches in the literature.
@@ 5.2 Theoretical Results
We now present our main theorem of this section, which states that we can achieve stability in the filtered search setting by choosing a sufficiently large distance penalty value. In this theoretical analysis, we assume that the attributes are drawn from some distribution such that the filter mismatch probability is at most some constant $p_{\max}$ where $0<p_{\max}<1$.
@@
We defer the full proof of this theorem to Appendix D. We note that Theorem 5.3 holds regardless of whether the underlying vector search problem (without constraints) is stable or unstable since our proof handles both cases directly. The latter case is particularly interesting as it implies that, under the assumptions of Theorem 5.3, the addition of filters can transform an unstable vector search problem into a stable one.
@@ 5.3 Experiments
In this section, we empirically validate Theorem 5.3 on a family synthetic vector search problems with 10,000 documents and 100 queries and varying dimensionality. We unit normalize all vectors so that the constant $\Delta\leq 2$ and set the parameter $p_{\max}$ to be $0.5$ which satisfies the conditions of the theorem. Thus, as predicted by our theory, a sufficient condition for stability based on these parameters is setting the penalty value $\alpha$ to be greater than $8$. For this experiment, we start with a base vector search problem that is unstable and we consider three different choices for the penalty value: in the first case we take the penalty to be 8.1, namely slightly above the minimum threshold for stability; secondly, we set the penalty value to be smaller than the threshold suggested by Theorem 5.3 and have it decay with the dimensionality; finally, we consider the case of a penalty value of 0. As suggested by the proof of Theorem 5.3, we want to engineer filter mismatches such that there is a negative correlation between the distance function and attribute matches. We simulate this phenomenon by setting filter matches and mismatches based on each document’s average distance to all queries; documents that are close to queries on average become filter mismatches while documents further away are matches.
@@ 6 Stability of Sparse Vector Search
We now turn our attention to the final vector search setting we study in this work: sparse embedding retrieval.
@@ 6.1 Theoretical Model
Restricting to nonnegative vectors entails no loss of generality. We provide a full proof of this claim in Lemma  E.1. We state our results for non-negative vectors for notational simplicity.
@@ 6.2 Theoretical Results
We defer the full proof of this theorem to Appendix E. This result is, to our knowledge, the first to prove sufficient conditions for stability in the context of sparse vector search, which is a particularly interesting setting given the high dimensionalities of learned sparse embeddings, such as SPLADE, used in practice.
The above theorem demonstrates that sparse vector search is stable whenever $X=(2-2\gamma)^{1/p}$ is less than $Y=\frac{\rho 2^{1/p}}{1-\pi}\left(\alpha^{1/p}-(1-\alpha)^{1/p}\right)$. These two expressions represent different geometric configurations for the vectors $q$ and $d$. The quantity $X$ is an upper bound on the expected distance when $q$ and $d$ overlap substantially in their head coordinates. In this case, the vectors partially align and the distance shrinks. In contrast, $Y$ represents the expected distance in the opposite configuration where $T_{q}$ and $T_{d}$ are disjoint. Here the vectors place their largest coordinates in opposite directions, and hence the distance grows.
Equality $X=Y$ corresponds to a single algebraic relationship between parameters $\alpha,\gamma$ and $\pi$ where the two geometric configurations happen to produce the exact same distance. Geometrically, this means that placing mass on overlapping head coordinates reduces the distance by precisely the same amount that shifting the mass into disjoint coordinates would increase it. This cancellation requires the parameters $(\alpha,\gamma,\pi)$ to satisfy a rigid relation and therefore occurs on a set of measure zero on the parameter space. Everywhere else for all practically relevant parameterizations, the two configurations remain geometrically distinct, which ensures that the distribution of the distances does not collapse in the limit, which we require for stability. Furthermore, in Appendix E.2, we empirically verify that the conditions of Theorem 6.4 hold true in practice on real SPLADE embeddings.
@@ 6.3 Stability of Synthetic Sparse Embeddings
Figure 3 shows the results of the experiments when we set $\alpha=0.83$, $\gamma=0.2$, and $\pi=0.5$ for the overlap probability. In particular, the CoI-only regime highlights that even strongly-concentrated embeddings exhibit distance collapse as the dimension grows when there is not enough overlap in their heads. These results confirm that neither concentration nor overlap of importance are sufficient conditions for stability on their own.
@@ 7 Conclusion
Appendix A Extended Related Work: Intrinsic Dimensionality
A closely related concept to stability is intrinsic dimensionality (ID) which measures the dimensionality of the subspace that a collection of vectors lie within. Intuitively, if a database of high-dimensional vectors lies within a much lower-dimensional manifold, then a near-neighbor search over these points is insulated from the curse of dimensionality effect. Thus, much like stability, ID provides a foundational theoretical tool for explaining why modern vector retrieval is capable of circumventing the curse of dimensionality. Consequently, ID has been well-studied with multiple proposed formal definitions. Perhaps the most influential definition of ID is that of (Karger and Ruhl, 2002) via a geomeric property called the expansion rate. Another popular definition heavily utilized in near-neighbor search algorithm design is the notion of doubling dimension proposed by (Gupta et al., 2003). We refer interested readers to the recent book of (Bruch, 2024) for the precise definitions. More recently, several works have measured the intrinsic dimensionality of modern neural embeddings and have reported that the ID is often much less than the ambient dimensionality. For instance, (Razzhigaev et al., 2024) study the intrinsic dimensionality of embeddings from transformer models (Vaswani et al., 2017) and find that the ID typically lies between 20 and 60 over the course of pretraining. In addition, (Gong et al., 2019) compute the intrinsic dimensionality of popular image model embeddings, finding that the ID in practice typically ranges from 12 to 19, well below the ambient dimensions of 128 to 512. Moreover, (Ansuini et al., 2019) go beyond embeddings and measure the ID of convolutional neural network (CNN) layers where they find the ID is an order of magnitude smaller than the ambient dimension.
Given that neural models consistently exhibit low intrinsic dimensionality, a natural question to ask is whether we still need to study why modern vector databases succeed in spite of the curse of dimensionality. On the contrary, we argue that intrinsic dimensionality alone is insufficient to explain the phenomena we see in practice. First and foremost, as we empirically measure in Section B, a dimensionality as low as 32, which is within the empirical ID measurements of real embeddings, is still high enough for the curse of dimensionality to diminish the search quality of approximate near neighbor search algorithms. This observation motivates the need to study stability explicitly. Secondly, modern embeddings produced by large language models and used in practice continue to grow in dimension into the thousands and tens of thousands of components. It is not clear if the intrinsic dimensionality will remain low as these embedding models scale and yet we observe they remain amenable to sub-linear vector search techniques. Moreover, stability analysis provides actionable guidance that ID alone cannot. For instance, our multi-vector stability theorem explains why ColBERT’s choice of Chamfer distance succeeds where average pooling might fail, a distinction invisible to ID measurements which only characterize the embedding space geometry, not the aggregation function. Similarly, our filtered search results suggest that practitioners can induce stability through sufficiently large penalties, even when the underlying embeddings are intrinsically high-dimensional.
Appendix B Practical Implications of Stability
In this section, we empirically demonstrate that stability is not merely a theoretical curiosity but also has substantial bearing on the quality of vector retrieval algorithms in practice. To our knowledge, this is the first empirical demonstration of the effects of instability on popular approximate ANN search algoritms. In Figure 4, we report the results of running two of the most popular ANN search techniques, namely the hierarchical navigable small worlds (HNSW) graph-based index (Malkov and Yashunin, 2018) and the inverted file (IVF) partitioning-based index (Douze et al., 2024), across stable and unstable datasets.
We fix a standard configuration of hyperparameters for each respective algorithm and compute the top-10 recall by taking the average of 1000 query vectors over a database of 1 million document vectors. To generate the unstable datasets at varying dimensionalities, we sample the value of each component independently from a univariate standard Gaussian distribution following the classic construction of (Beyer et al., 1999). To construct the stable datasets, we generate vectors such that they belong to one of five distinct clusters. For consistency, we fix the search index hyperparameters across all experiments; for HNSW, we set ef-construction and ef-search to be $200$ and $\texttt{M}=16$; for IVF, we set the number of clusters to $\texttt{nlist}=\sqrt{n}$, where $n$ is the number of database elements and the number of probes to $\texttt{nprobes}=\frac{\sqrt{n}}{4}$.
From Figure 4, we see that these approximate near-neighbor search algorithms severely suffer in recall over unstable databases since the lack of contrast between points renders sublinear algorithms that avoid a brute force scan ineffective. We also see that the impact of instability worsens in higher dimensions. These results underscore the importance of stability in practice.
Appendix C Stability of Multi-Vector Search
C.1 Proofs
@@
Theorem 4.5 For sufficiently large databases, a multi-vector search instance is stable if and only if
@@
This result follows directly from the prior proofs of (Beyer et al., 1999) and (Durrant and Kabán, 2009) since the original proofs require only that the distance metric is non-negative. We provide a full, self-contained argument here for completeness.
As a matter of notation, let $\mathrm{Agg}(Q,D)$ be the random variable for the aggregated distance in an $m$-dimensional space. Let its expected value be $\mathbb{E}[\mathrm{Agg}(Q,D)]$ and its variance be $\text{Var}[\mathrm{Agg(Q,D)}]$. The database produces $n$ samples of this random variable. Let $\mathrm{DMIN}=\min_{D\in\mathcal{D}}\mathrm{Agg}(Q,D)$ and $\mathrm{DMAX}=\max_{D\in\mathcal{D}}\mathrm{Agg}(Q,D)$.
$\implies$ This direction follows the argument from (Beyer et al., 1999). We will show that if $\lim_{m\to\infty}\frac{\text{Var}[\mathrm{Agg}(Q,D)]}{\mathbb{E}[\mathrm{Agg}(Q,D)]^{2}}=0$, then the multi-vector search problem is unstable. Define a normalized random variable $Y=\frac{\mathrm{Agg}(Q,D)}{\mathbb{E}[\mathrm{Agg}(Q,D)]}$. Its expectation is $\mathbb{E}[Y]=1$ and its variance is $\text{Var}[Y]=\frac{\text{Var}[\mathrm{Agg}(Q,D)]}{\mathbb{E}[\mathrm{Agg}(Q,D)]^{2}}$ The condition is thus equivalent to $\lim_{m\to\infty}\text{Var}[Y]=0$. Since $\lim_{m\to\infty}\mathbb{E}[Y]=1$ and $\lim_{m\to\infty}\text{Var}[Y]=0$, the random variable $Y$ converges in probability to 1. Let $\{Y_{1},\dots,Y_{n}\}$ be the normalized distances for the $n$ items in the database. Since $\min()$ and $\max()$ are continuous functions, Slutsky’s Theorem implies that $\min Y_{i}\to_{p}1$ and $\max Y_{i}\to_{p}1$. Since both the numerator and denominator converge in probability to 1, their ratio also converges in probability to 1. By definition, $\frac{\mathrm{DMAX}}{\mathrm{DMIN}}\to_{p}1$ means that for any $\epsilon>0$, we have $\lim_{m\to\infty}\Pr\left[\left|\frac{\mathrm{DMAX}}{\mathrm{DMIN}}-1\right|\leq\epsilon\right]=1$. Since $\mathrm{DMAX}\geq\mathrm{DMIN}$, this is equivalent to $\lim_{m\to\infty}\Pr\left[\mathrm{DMAX}\leq(1+\epsilon)\mathrm{DMIN}\right]=1$. This is precisely the definition of an unstable search problem as stated in Definition 4.4.
@@
$\frac{\mathrm{DMIN}}{\mathrm{DMAX}}\leq\frac{\mathrm{Agg}(Q,D_{i})}{\mathbb{E}[\mathrm{Agg}(Q,D_{i})]}\leq\frac{\mathrm{DMAX}}{\mathrm{DMIN}}$
From the instability assumption, we know both the lower bound $\frac{\mathrm{DMIN}}{\mathrm{DMAX}}$ and the upper bound $\frac{\mathrm{DMAX}}{\mathrm{DMIN}}$ converge in probability to 1 By the Squeeze Theorem, the random variable $Y$ must also converge in probability to 1. A random variable that converges in probability to a constant must have a variance that converges to zero. Therefore:
$\lim_{m\to\infty}\text{Var}[Y]=\lim_{m\to\infty}\frac{\text{Var}[\mathrm{Agg}(Q,D)]}{\mathbb{E}[\mathrm{Agg}(Q,D)]^{2}}=0$
@@
$\frac{\mathrm{DMAX}^{\prime}}{\mathrm{DMIN}^{\prime}}=\frac{\max_{k}\min_{d\in D_{k}}\delta(q,d)}{\mathrm{DMIN}}\geq\frac{1}{c}\frac{\mathrm{DMAX}}{\mathrm{DMIN}}>\frac{1}{c}(c+\epsilon)>1.$
for some constant $\epsilon>0$ by $c$-strong stability. Thus, it follows that $\frac{\mathrm{DMAX}^{\prime}}{\mathrm{DMIN}^{\prime}}\not\to_{p}1$, completing the proof. ∎
@@
Lemma C.1 (Dan’s Favorite Inequality (Spielman, 2018)).
@@
where the last two inequalities follow from applying Lemma C.1 and Lemma 4.8 respectively. This completes the proof.
Remark C.2.
Stability follows from applying Dan’s Favorite Inequality (Lemma C.1) to reduce the general multi-vector search problem to multi-vector search over singleton query sets, namely the setting in Lemma 4.8, at the cost of an $\frac{1}{k}$ overhead factor. The $1/k$ factor suggests that stability weakens as the query set size grows, which is also consistent with empirical observations.
C.2 Empirical Validation of Multi-Vector Theorem Assumptions
Theorem 4.9 relies on three assumptions to prove the stability of multi-vector search under Chamfer distance: $c$-strong stability, non-degenerate document sets, and a non-negative bound on the sum of the covariances of the individual Chamfer distance terms. In this section, we experimentally demonstrate that all three of these assumptions are satisfied in practice, which underscores the practical utility of our result. Specifically, we use ColBERT embeddings (Khattab and Zaharia, 2020) to construct query and document vector sets for five standard English-based information retrieval datasets from the BEIR benchmark (Thakur et al., ). For each dataset, we empirically estimate the strong stability constant $c$ as the smallest value of the ratio between the maximum distance and the minimum distance across all queries in the induced single-vector search instance. We then use this value of $c$ to check the proportion of document sets that satisfy the non-degeneracy condition. Finally, we then compute the sum of the covariances of the near-neighbor distances of individual query vectors in a given query set and verify that the sum is non-negative. This non-negative covariance assumption is natural in practice since the tokens in a query set tend to be topically related. We summarize our empirical measurements in Table 1 where we see that all three theorem conditions are comfortably satisfied in practice.
Dataset Stability Constant ($c$) Non-Degeneracy Pass Rate (%) Cov. Bound Pass Rate (%)
Natural Questions 2.36 100.0 100.0
HotpotQA 2.12 100.0 100.0
TREC-COVID 2.59 100.0 100.0
NFCorpus 1.91 100.0 100.0
Requirement >1>1 =100=100 =100=100
C.3 Stability of ColBERT Embeddings
Appendix D Stability of Filtered Vector Search